# Bridging the Gap between Discrete Agent Strategies in Game Theory and Continuous Motion Planning in Dynamic Environments

### v1

@@ IV-B Offline Phase: Motion Planning and Policy Synthesis
We use a sampling-based motion planner. First, the planner samples local goals for the vehicle in a lattice pattern, as shown in Figure 4. Then, dynamically feasible trajectories are generated [27] that take the vehicle from the current state to the sampled goal state. For each sampled candidate trajectory, we calculate a few cost functions that correspond to geometric properties (e.g. maximum curvature, deviation from the race line, etc.). The trajectory for the agent to follow is the candidate trajectory that has the lowest weighted sum of all cost function values. We parameterize the motion planner with the weight vector. More details on the cost functions used can be found in Appendix A. After selecting a trajectory, the vehicle control input is generated using Pure Pursuit [28]. We can see the different behavior of the agent induced by having different cost weights in Figure 4. The planner on the left weighs the penalty on deviating from a predetermined raceline less, thus generating a path that keeps a higher velocity around the opponent vehicle. The planner on the right weights this penalty more, thus generating a path that slows down and follows the opponent in front, while keeping a trajectory closely matching the raceline.
@@ IV-C Online: Strategy Opitmization with Approximated CFR
Before online strategy optimization, we need to first train the regret approximator $\mathcal{R}_{\operatorname{approx}}(h,e_{k})$ that estimates the counterfactual regret for each action given a history at the final iteration of CFR. Training data are collected by playing full extensive games represented as game trees. First, we allow each agent to take $m$ actions in total in a game. The initial starting policies for both agents in the root node are chosen as random policies from $\Pi_{p}$. Then, we traverse every single branch on the game tree by taking all combinations of action at each node for both agents. At the terminal nodes of the tree, the final utilities are calculated for each game, and the corresponding counterfactual regret is also calculated for every action for every corresponding history present on the tree. The total number of games played between two agents is $N_{\text{init}}^{2}(2{|\mathcal{G}|})^{2m}$, where $N_{\text{init}}$ is the number of initial starting policies for each agent. For more details on how subsets of policies used are selected, and the size of the training dataset, please refer to Appendix B. The number of games played, even considering the relatively small number of $N_{\text{init}}$ and $m$, is still massive. This further strengthens the need for an approximator for counterfactual regret.
@@ VI Conclusion
Appendix A Motion Planner Cost Functions
We define multiple cost functions to evaluate the quality of each trajectory for geometric properties and velocity profiles. We use the following set of cost functions $f_{i}\in\mathcal{F}$:
  • $f_{\text{mc}}=\max_{s=0}^{s_{f}}\{\kappa_{s}\}$: maximum curvature on the trajectory
  • $f_{\text{al}}=$ arc length $s_{f}$ of the trajectory
  • $f_{\text{hys}}=\left|\left|\mathcal{T}_{t}-\mathcal{T}^{*}_{t-1}\right|\right|_{2}$ hysteresis loss that measures similarity to the previously selected trajectory. Calculated as the Euclidean distance between the two trajectories. Note that the previously selected trajectory is shifted forward to compensate for the vehicle’s motion between time steps.
  • $f_{\text{do}}$ distance to the optimal race line measured by lateral deviation. For every single position on the trajectory, a corresponding nearest point is found on the optimal raceline. Then the lateral is calculated in the Frenet coordinate frame.
  • $f_{\text{co}}$ fixed collision cost with the opponent discounted by relative speed to the opponent at each time step.
  • $f_{\text{v1}}$ velocity cost that encourages higher speed.
  • $f_{\text{v2}}$ velocity cost that penalizes co-occurrence of high speed and high curvature.
The motion planner selects the target trajectory by:
$\mathcal{T}^{*}=\underset{\Theta}{\operatorname{\text{argmin}}}\sum_{\theta_{i}\in\Theta,f_{i}\in\mathcal{F}}\theta_{i}f_{i}$ (11)
Therefore, each agent can be parameterized by a weight vector:
$\Theta=\left[~{}\theta_{mc},~{}\theta_{al},~{}\theta_{\text{hys}},~{}\theta_{do},~{}\theta_{co},~{}\theta_{v1},~{}\theta_{v2}\right]$
In addition, trajectories in collision with the environment are assigned infinite costs.
Appendix B Regret Approximator Training
MO-CMA-ES results in two sets of policies: all sampled policies $\Pi_{\text{all}}$ and policies on the Pareto front $\Pi$. Due to $\Pi$ being sparse, we upsample it into a near-optimal set using policies within a certain distance to each point in $\Pi$. We use Determinantal Point Processes (DPP) [31] to create subsets used during both the offline and online portion of the method. Figure 7 shows how all sets of policies compare in the Policy Characteristics Space.
Since the number of discrete actions is determined by the number of policy characteristic functions $g_{c}$s chosen, we know the number of admissible actions is $\left|\mathcal{E}\right|=2{|\mathcal{G}|}$. With $m$ as the number of allowed actions for each agent, each game tree has depth $2m$, and the total number of branches $(2{|\mathcal{G}|})^{2m}$. If we select $N_{\text{init}}$ initial starting policies for each agent, there will be $N_{\text{init}}^{2}$ number of game trees with different root nodes. Thus, in total, the number of data points collected to train the approximator is $N_{\text{init}}^{2}(2{|\mathcal{G}|})^{2m}$. Next, we describe how to construct the games played for data collection. First, two subsets of the near-optimal set $\Pi_{\text{DPP1}}$ and $\Pi_{\text{DPP2}}$, where $\Pi_{\text{DPP1}}\cap\Pi_{\text{DPP2}}=\varnothing$, and $\left|\Pi_{\text{DPP1}}\right|=\left|\Pi_{\text{DPP2}}\right|$. We construct game trees using each pair in the Cartesian product $\Pi_{\text{DPP1}}\times\Pi_{\text{DPP2}}$ as the initial starting points in the Policy Characteristic Space of the two agents. We then play through the entire game tree and collect counterfactual regret for each history and action pair for training. We used a multilayer perceptron (MLP) with one hidden layer of size 2048 and ReLU activation as the approximator. The network is trained using L1 loss on the prediction targets, and optimized with Adam with adaptive learning rate (starting at 0.005, reduce when validation loss plateaus for ten steps) for 2000 epochs with a batch size of 1024.
In our experiments, we choose game tree depths of 6, dimension of the Policy Characteristics Space $|\mathcal{G}|=2$, and the size of starting set of policies for each agent $N_{\text{init}}=20$. The total number of games played during data collection is $163840020^{2}\times 4^{6}=$ 1638400.
Figure 7: Subsets created from offline policy synthesis to be used in the online portion of proposed method.
Appendix C Action discretization comparisons
TABLE II: Success rate and elapsed times of different agents finishing two laps in a single agent setting.
Agent Success Rate Avg. Elapsed Time (s)
On Seen Map
PPO-continuous 0.0 N/A
PPO-discrete 0.3 $63.782\pm 0.225$
Ours (w/o CFR) 1.0 $48.533\pm 1.398$
On Unseen Map
PPO-continuous 0.0 N/A
PPO-discrete 0.25 $67.292\pm 0.156$
Ours (w/o CFR) 1.0 $50.814\pm 1.343$
We compare three agents in this experiment. The first is PPO [32] with continuous actions on both steering and throttle. The second is PPO with discretized actions: turn left, turn right, and stay straight. The third is our proposed approach without the CFR updates. The agents under both PPO settings are rewarded by a small keep alive reward for every time step the car is not in collision and a large terminal reward when finishing two laps under the time limit. They are also penalized by a value scaled with lap time. The PPO policies use the range measurement vector from the LiDAR scan as the observation. During training, the PPO agent with discrete actions converges to a policy that can finish two laps. In comparison, the PPO agent with continuous actions does not. We set up the experiments as follows. For PPO agents, the same trained agent is used in all rollouts, and the starting positions of the agents are slightly different in each rollout. For agents using the proposed approach, we again sample a subset of the near-optimal set using DPP. The number of DPP samples here matches the number of random starts for the PPO agents. We record the success rate and lap times over 20 trials for each agent both on the seen map and on the unseen map.
The recorded results are shown in Table II. First, we compare the two PPO agents to see the effect of using discretized actions. Our experiment confirms that agents with discretized actions are easier to train. At the end of training, the continuous PPO agent, although turning in the correct direction when encountering corners, was unable to complete the entire lap. Then we compare the discrete PPO with our proposed method. Although able to complete the two laps during training, simply changing the starting position by a small amount drops the success rate to only 30% even when evaluations are performed on the same map the agent saw during training. When moved to an unseen map, the success rate drops further to only 25%. In comparison, the proposed method maintains a perfect success rate throughout the evaluations, even when moved to an unseen map. Although not the main objective of this experiment, we see that the proposed method was able to achieve a lower lap time across the board. Although PPO agents are awarded for having shorter lap times, they cannot compete with the proposed method.
Figure 8: Episodic returns when training PPO policies. Moving average of 10 steps shown.
RL agents used for the comparison are trained using continuous and discrete PPO in CleanRL [33]. We show the episodic returns for each agent in 8. The training is capped at 1000000 total time steps, with 7000 steps between each policy update. The learning rate is annealed and starts at 3e-4 for continuous PPO and 2.5e-4 for discrete PPO. The discount factor for both is set at 0.99. The $\lambda$ for the general advantage estimation for both is set at 0.95. The epochs to update is 10 for the continuous policy and 4 for the discrete policy. The surrogate clipping coefficient is 0.2 for both policies. The entropy coefficient is 0 for the continuous policy and 0.01 for the discrete policy. The coefficient for the value function is 0.5 for both policies. The maximum norm for gradient clipping is 0.5 for both policies. The continuous action is sampled from a learned Gaussian with control input limits. The discrete action is steering with 0.3, $-0.3$, and 0.0 steering angles, all with a velocity of 5.0 m/s. Agents get a reward of 0.01 for every time step not in collision, a reward of 40 for reaching the end of two laps without collision. The gym environment is set with a timeout to avoid infinite episodes.

### v2

# Bridging the Gap between Discrete Agent Strategies in Game Theory and Continuous Motion Planning for Dynamic Environments

@@ IV-B Offline Phase: Motion Planning and Policy Synthesis
We use a sampling-based motion planner. First, the planner samples local goals for the vehicle in a lattice pattern, as shown in Figure 4. Then, dynamically feasible trajectories are generated [27] that take the vehicle from the current state to the sampled goal state. For each sampled candidate trajectory, we calculate a few cost functions that correspond to geometric properties (e.g. maximum curvature, deviation from the race line, etc.). The trajectory for the agent to follow is the candidate trajectory that has the lowest weighted sum of all cost function values. We parameterize the motion planner with the weight vector. More details on the cost functions used can be found in Appendix LABEL:app:costs. After selecting a trajectory, the vehicle control input is generated using Pure Pursuit [28]. We can see the different behavior of the agent induced by having different cost weights in Figure 4. The planner on the left weighs the penalty on deviating from a predetermined raceline less, thus generating a path that keeps a higher velocity around the opponent vehicle. The planner on the right weights this penalty more, thus generating a path that slows down and follows the opponent in front, while keeping a trajectory closely matching the raceline.
@@ IV-C Online: Strategy Opitmization with Approximated CFR
Before online strategy optimization, we need to first train the regret approximator $\mathcal{R}_{\operatorname{approx}}(h,e_{k})$ that estimates the counterfactual regret for each action given a history at the final iteration of CFR. Training data are collected by playing full extensive games represented as game trees. First, we allow each agent to take $m$ actions in total in a game. The initial starting policies for both agents in the root node are chosen as random policies from $\Pi_{p}$. Then, we traverse every single branch on the game tree by taking all combinations of action at each node for both agents. At the terminal nodes of the tree, the final utilities are calculated for each game, and the corresponding counterfactual regret is also calculated for every action for every corresponding history present on the tree. The total number of games played between two agents is $N_{\text{init}}^{2}(2{|\mathcal{G}|})^{2m}$, where $N_{\text{init}}$ is the number of initial starting policies for each agent. For more details on how subsets of policies used are selected, and the size of the training dataset, please refer to Appendix LABEL:app:nn. The number of games played, even considering the relatively small number of $N_{\text{init}}$ and $m$, is still massive. This further strengthens the need for an approximator for counterfactual regret.